# Cyberattack Detection in Virtualized Microgrids Using LightGBM and Knowledge-Distilled Classifiers

### v1

@@ III-C Secondary Control Interface for Attack Injection
III-C1 Ramp Attack
@@
III-C2 Slow Ramp Attack
@@ III-C3 Sinusoidal Attack
III-C4 Additive Bias Attack
$\xi_{i}^{\text{mal}}(t)$ $=\xi_{i}(t)+b,$ (13)
$\zeta_{i}^{\text{mal}}(t)$ $=\zeta_{i}(t)+b.$ (14)
@@ III-E Noise Representation and Robustness Modeling
Although cyberattacks introduce significant disturbances into the microgrid secondary control layer, inverter-based microgrids naturally operate under several sources of physical and numerical noise even in the absence of malicious behavior. Because these disturbances influence voltage, current, power, and frequency measurements, they directly affect the data used to train machine learning models. To ensure that the learning algorithms developed in this work generalize to real-world operating conditions, the virtual microgrid was simulated using a switching-level model with a fixed $2~\mu\text{s}$ step size. As depicted in Fig. 5, this configuration inherently produces structured noise and variability consistent with practical microgrid deployments, consistent with the recommendations in [39] and the cyber-physical control behavior described in the reference microgrid architecture.
@@ V Results
This section evaluates the proposed cyberattack-detection framework, covering binary detection, multiclass attack classification, feature-importance analysis, knowledge-distillation results, and real-time prediction performance. All experiments were performed on the normalized, downsampled dataset produced from the high-resolution simulations described in Section IV.
V-A Binary Attack Detection Performance
@@
V-B Multiclass Attack-Type Classification
@@ V-B1 Test-Set Performance
V-B2 Confusion Matrix Interpretation
@@ V-C LightGBM Hyperparameters
V-D Statistical Robustness and Ablations
@@
V-E Feature-Importance Analysis
@@
V-F Knowledge Distillation Evaluation
@@
As shown in Fig. 8, the student achieved similar results as the teacher model.
@@
The knowledge-distilled student preserves diagnostic performance while reducing inference time by approximately 73% relative to the teacher model. This makes the detector suitable for microgrid edge controllers with limited processing power.
@@
V-G Real-Time Prediction Behavior
@@
As shown in Fig. 9, the trained multiclass LightGBM model was then used to make a prediction for each selected sample, one point at a time. Each predicted label was directly compared with its corresponding ground-truth label in the test set, and all ten predictions matched perfectly. This confirms that the classifier generalizes to unseen temporal slices and decision boundaries remain stable under noise and switching ripple. The system behaves reliably for point-wise real-time inference.
@@ VI Computational Efficiency and Deployment Considerations
Effective cyberattack detection in microgrids must be both accurate and fast [22]. The detection algorithm also needs to run with low computational cost and meet strict real-time requirements [31]. This section evaluates the computational efficiency of the LightGBM-based detection framework, including model size, inference speed, CPU-only performance, and the improvements gained through knowledge distillation. All runtime measurements correspond to the same experimental setup used in Section V.
@@ VI-A Hardware Environment
All inference-time and model-size measurements were performed on a virtualized Google Colab CPU instance. The system uses a GenuineIntel Intel Xeon @ 2.20 GHz processor with the following configuration:
  • 1 physical core, 2 threads
  • x86_64 architecture
  • KVM hypervisor virtualization
  • 55 MB L3 cache
  • AVX, AVX2, FMA extensions enabled
This environment represents a modest CPU-only computing platform, allowing fair comparison with embedded-class processors that lack GPU acceleration.
@@ VI-C Inference-Time Benchmarking
A major benefit of the proposed framework is its compatibility with CPU-only environments. Unlike CNNs, LSTMs, or transformer-based detectors that need many floating-point operations and often rely on GPU acceleration, LightGBM uses simple operations such as threshold comparisons, integer additions, and shallow tree traversal. These operations are extremely efficient on embedded processors.
In contrast, CNN-based smart grid detectors [18, 42] rely on millions of convolution operations. LSTM-based microgrid models 11 require large matrix multiplications and memory-intensive time-state storage. The proposed LightGBM+KD framework avoids these bottlenecks entirely.
@@ VI-F4 Extendable taxonomy
New cyberattack modes can be integrated simply by retraining with additional labeled datasets.
@@ VI-F5 Compatibility with federated learning
Techniques similar to [25] can be applied so that decentralized microgrids collaboratively improve the detection model without sharing raw data.
@@ VI-G Implication for Future Systems
Firstly, multi-DER distribution grids, which increasingly rely on inverter-interfaced resources, can benefit from having lightweight anomaly detectors co-located with local controllers [30]. The capability to run efficiently on CPU-only hardware makes on-device inference feasible even at high sampling rates. Furthermore, inverter-dominated AC and DC microgrids can integrate the framework at the secondary-control or tertiary-control layers without modifying existing control hardware [14]. This positions the model as a practical cybersecurity safeguard to accompany power-sharing and voltage-restoration mechanisms. Thirdly, the method aligns naturally with edge-to-cloud architectures. A local LightGBM model can run at the edge (e.g., on DG controllers or protection relays), while aggregated attack indicators or distilled student models may be shared with supervisory systems or cloud-hosted analytics platforms [40].
Finally, the framework is compatible with hybrid AC/DC microgrid environments, which require fast decision-making across heterogeneous converter topologies. Overall, the combination of high accuracy, low inference latency, and modest computational requirements positions the proposed system as a strong candidate for next-generation resilient smart grids. It supports real-time cyberattack detection across many types of microgrid setups and can easily scale to new distributed energy systems.
@@ VII Conclusion
This work presented a complete cyber-physical framework for generating, labeling, and classifying secondary-control cyberattacks in inverter-based microgrids. A high-fidelity virtual microgrid was constructed in MATLAB/Simulink using distributed droop-controlled DGs, consensus-based secondary controllers, and structured cyberattack injection pathways. The simulation environment recorded detailed physical and cyber behavior at a very high time resolution of $2~\mu\text{s}$. This made it possible to generate realistic and well-labeled datasets that accurately represent both normal operation and different types of cyberattacks.
A complete machine learning pipeline was built to process the generated datasets. After applying incremental normalization and balanced downsampling, we trained binary and multiclass LightGBM models to detect and classify six representative secondary-control attack types. The multiclass model reached 99.7% accuracy, showing that the attack types can be recognized correctly by the classifier. Knowledge distillation was then used to compress the model, creating a smaller student classifier that kept the same accuracy while reducing the inference time by about 73%.
The results show that the tree-boosted models provide a strong mix of accuracy, reliability, and fast computation. Their compact structure enables practical real-time deployment across a wide range of microgrid controllers. Furthermore, the structured dataset and modeling approach provide a repeatable blueprint for future cyber-resilient microgrid studies. Future work will extend the framework to include (i) additional DER types and converter topologies, (ii) communication-layer adversaries with time-varying topologies, (iii) reinforcement-learning-based adaptive controllers, and (iv) federated or privacy-preserving architectures to enable secure multi-microgrid coordination.
Overall, this study shows that using high-resolution cyber-physical simulations together with lightweight gradient-boosting models and knowledge distillation provides a practical and effective way to detect cyberattacks in real time within modern microgrids.

### v2

@@ III-C Secondary Control Interface for Attack Injection
III-C1 Additive Bias Attack
$\xi_{i}^{\text{mal}}(t)$ $=\xi_{i}(t)+b,$ (9)
$\zeta_{i}^{\text{mal}}(t)$ $=\zeta_{i}(t)+b.$ (10)
III-C2 Ramp Attack
@@
III-C3 Slow Ramp Attack
@@ III-E Noise Representation and Robustness Modeling
Although cyberattacks disturb the secondary-control layer, inverter-based microgrids also contain natural physical and numerical noise even during normal operation. This noise comes mainly from high-frequency inverter switching, digital sampling, and discretized control updates. These effects introduce small ripples and variability into measured voltage, current, power, and frequency signals. To capture these realistic disturbances, the virtual microgrid was simulated using a switching-level inverter model with a fixed $2~\mu\text{s}$ step size. As shown in Fig. 5, this produces switching ripple at the electrical level, staircase-like sampled measurements at the control level, and discrete-time communication of consensus variables in the secondary controller. Together, these effects create structured noise and variability similar to what is observed in practical inverter-based microgrids [39]. This ensures that the machine-learning models are trained on data that reflect real operating conditions rather than idealized, noise-free signals.
@@ V Results
This section evaluates the proposed cyberattack-detection framework. It first examines how secondary-control attacks appear in physical measurements, and then reports the performance of the machine-learning models. The results cover binary attack detection, multiclass attack classification, feature-importance analysis, knowledge-distillation outcomes, and real-time prediction behavior. All experiments were performed on the normalized, downsampled dataset produced from the high-resolution simulations described in Section IV.
V-A Physical Signatures of Secondary-Control Attacks
To show how secondary-control attacks appear in physical measurements, Fig. 6 shows the time-domain response of DG1 under an additive-bias attack. The figure compares normal and attacked values of bus voltage $V_{1}(t)$, active power $P_{\mathrm{DG1}}(t)$, reactive power $Q_{\mathrm{DG1}}(t)$, and frequency $f_{\mathrm{DG1}}(t)$.
Figure 6: Representative DG1 time-domain response under an additive-bias attack injected at $t_{a}=0.70$ s. The attack perturbs the secondary-control consensus signals, producing measurable deviations in (a) bus voltage $V_{1}(t)$, (b) active power $P_{\mathrm{DG1}}(t)$, (c) reactive power $Q_{\mathrm{DG1}}(t)$, and (d) frequency estimate $f_{\mathrm{DG1}}(t)$ compared with normal operation.
Even over a short time window, the additive attack causes visible changes in all four variables. The attacked curves slowly move away from the normal ones, showing that a bias added at the secondary-control layer passes through the droop controller and inner control loops and appears in the electrical outputs of DG1. Changes in $P_{\mathrm{DG1}}(t)$ and $Q_{\mathrm{DG1}}(t)$ show that power sharing is altered, while deviations in $V_{1}(t)$ and $f_{\mathrm{DG1}}(t)$ show weaker voltage and frequency regulation. Although the change in any single variable may look small, the combined behavior of $(V,P,Q,f)$ forms a clear pattern that can be used for attack detection.
Figure 7: Representative DG1 time-domain responses under different secondary-control cyberattacks compared with normal operation. (a) Ramp attack: bus voltage $V_{1}(t)$. (b) Ramp attack: active power $P_{\mathrm{DG1}}(t)$. (c) Slow-ramp attack: reactive power $Q_{\mathrm{DG1}}(t)$. (d) Sinusoidal attack: reactive power $Q_{\mathrm{DG1}}(t)$. (e) Coordinated stealth attack: active power $P_{\mathrm{DG1}}(t)$. (f) Denial-of-service (DoS) attack: bus voltage $V_{1}(t)$. Each panel compares the normal (blue) and attacked (orange) trajectories.
Fig. 7 compares the effects of several secondary-control attacks on DG1.
Ramp attacks cause growing distortion in $V_{1}(t)$ and a steady drift in $P_{\mathrm{DG1}}(t)$, showing gradually changing power sharing. Slow-ramp attacks create a small but persistent bias in $Q_{\mathrm{DG1}}(t)$. Sinusoidal attacks add oscillations and a clear offset to $Q_{\mathrm{DG1}}(t)$. Coordinated stealth attacks cause only small changes in $P_{\mathrm{DG1}}(t)$, and the attacked curve stays close to the normal one. This shows how low-amplitude coordinated disturbances can look like normal operation while still degrading performance. DoS attacks produce larger oscillations and peaks in $V_{1}(t)$ because frozen or delayed consensus updates weaken voltage regulation.
Overall, different secondary-control attacks leave different physical patterns in $(V,P,Q)$. Ramp and DoS attacks create clearly visible changes, while slow-ramp and stealth attacks cause small but consistent shifts. These combined changes across multiple variables provide useful features for the LightGBM classifier to detect attacks and identify their type.
V-B Binary Attack Detection Performance
@@
V-C Multiclass Attack-Type Classification
@@ V-C1 Test-Set Performance
V-C2 Confusion Matrix Interpretation
@@ V-D LightGBM Hyperparameters
V-E Statistical Robustness and Ablations
@@
V-F Feature-Importance Analysis
@@
V-G Knowledge Distillation Evaluation
@@
As shown in Fig. 10, the student achieved similar results as the teacher model. The knowledge-distilled student preserves diagnostic performance while reducing inference time by approximately 73% relative to the teacher model. This makes the detector suitable for microgrid edge controllers with limited processing power.
@@
Time-Domain Prediction Consistency
In addition to aggregate metrics, we examine whether the distilled student preserves the teacher’s decision behavior over time. Fig. 11 plots the ground-truth class sequence and the corresponding teacher and student predictions over a short time-ordered window. A small vertical offset is applied only for visualization, so the three trajectories remain distinguishable. This result supports using the student model for real-time deployment, since it reproduces the teacher’s switching behavior with minimal deviation while providing significantly lower inference time.
Figure 11: Ground-truth, teacher, and student (KD) discrete class trajectories over a short time window. Blue circles indicate the true class, red squares show the teacher (LightGBM) predictions, and yellow triangles show the student (KD) predictions. A small vertical offset is applied for visual separation. The student closely follows the teacher’s class transitions and matches the ground truth in nearly all samples, indicating that KD preserves temporal decision behavior.
V-H Real-Time Prediction Behavior
@@
As shown in Fig. 12, the trained multiclass LightGBM model was used to predict each selected sample one at a time. Each prediction was compared with its ground-truth label, and all ten matched correctly. This shows that the classifier works well on unseen data and remains stable under noise and switching ripple. The system is reliable for point-wise real-time inference.
@@ VI Computational Efficiency and Deployment Considerations
Effective cyberattack detection in microgrids must be both accurate and fast [22]. The detection algorithm must also run with low computational cost and meet strict real-time requirements [31]. This section evaluates the computational efficiency of the proposed LightGBM-based framework, including model size, inference speed, CPU-only performance, and the benefits of knowledge distillation. All runtime measurements use the same experimental setup as in Section V.
@@ VI-A Hardware Environment
All inference-time and model-size measurements were performed on a virtualized Google Colab CPU instance. The system used an Intel Xeon processor (GenuineIntel) running at 2.20 GHz with one physical core and two hardware threads, based on the x86_64 architecture and virtualized using a KVM hypervisor. The processor provided a 55 MB L3 cache and supported AVX, AVX2, and FMA instruction sets. This environment represents a modest CPU-only computing platform, allowing fair comparison with embedded-class processors that lack GPU acceleration.
@@ VI-C Inference-Time Benchmarking
A major benefit of the proposed framework is its compatibility with CPU-only environments. Unlike CNNs, LSTMs, or transformer-based detectors that need many floating-point operations and often rely on GPU acceleration, LightGBM uses simple operations like threshold comparisons, integer additions, and shallow tree traversal. These operations are very efficient on embedded processors. In contrast, CNN-based smart grid detectors [18, 42] use millions of convolution operations, which are slow and computationally heavy. LSTM-based microgrid models need large matrix calculations and store many internal states in memory. The proposed LightGBM+KD framework avoids these heavy computations and memory costs.
@@ VI-F4 Extendable taxonomy
New cyberattack modes can be integrated by retraining with additional labeled datasets.
@@ VI-F5 Compatibility with federated learning
Techniques similar to [25] can be applied so that decentralized microgrids improve the detection model without sharing raw data.
@@ VI-G Implication for Future Systems
Firstly, multi-DER distribution grids, which increasingly rely on inverter-interfaced resources, can benefit from having lightweight anomaly detectors co-located with local controllers [30]. Because the model runs well on CPU-only hardware, it can make real-time decisions even at high sampling rates. Furthermore, inverter-based AC and DC microgrids can add the framework at the secondary or tertiary control level without changing existing control hardware [14]. This makes the model a practical cybersecurity add-on for power-sharing and voltage-restoration control. Thirdly, the method aligns naturally with edge-to-cloud architectures. A local LightGBM model can run at the edge (e.g., on DG controllers or protection relays), while aggregated attack indicators or distilled student models may be shared with supervisory systems or cloud-hosted analytics platforms [40].
Finally, the framework is compatible with hybrid AC/DC microgrids that require fast decision-making across heterogeneous converter topologies. The combination of high accuracy, low inference latency, and modest computational requirements makes the proposed system a strong candidate for next-generation resilient smart grids. It enables real-time cyberattack detection across a wide range of microgrid configurations and can be readily extended to new distributed energy systems.
@@ VII Conclusion
This work presented a complete cyber-physical framework for generating, labeling, and classifying secondary-control cyberattacks in inverter-based microgrids. A high-fidelity virtual microgrid was built in MATLAB/Simulink using distributed droop-controlled DGs, consensus-based secondary controllers, and structured cyberattack injection pathways. The simulation environment recorded detailed physical and cyber behavior at a very high time resolution of $2~\mu\text{s}$. This made it possible to generate realistic and well-labeled datasets that represent both normal operation and different types of cyberattacks. A complete ML pipeline was built to process the generated datasets. After applying incremental normalization and balanced downsampling, we trained binary and multiclass LightGBM models to detect and classify six representative secondary-control attack types. The multiclass model reached 99.7% accuracy, showing that the attack types can be recognized correctly by the classifier. KD was then used to compress the model, creating a smaller student classifier that kept the same accuracy while reducing the inference time by about 73%.
The results show that the tree-boosted models provide a strong mix of accuracy, reliability, and fast computation. Their compact structure enables practical real-time deployment across a wide range of microgrid controllers. Furthermore, the structured dataset and modeling approach provide a repeatable blueprint for future cyber-resilient microgrid studies. Future work will extend the framework to include (i) additional DER types and converter topologies, (ii) communication-layer adversaries with time-varying topologies, (iii) Reinforcement Learning-based adaptive controllers, and (iv) federated or privacy-preserving architectures to enable secure multi-microgrid coordination. Overall, this study shows that using high-resolution cyber-physical simulations together with lightGBM models and KD provides a practical and effective way to detect cyberattacks in real time within modern microgrids.